Two charged objects are connected to a spring with
Answer #1

40) The force on the second are

kx - kq^2/R^2 = 0 ;

500 * (1-.5) = k q^2/R2

q = sqrt (250/9*109)

q = 167 uC

41 ) writing the similar equation

K ( .5 - R ) = k Q^2/R^2

250 R^2 - 500 R^3 = 9*109 * 9*10-10

250 R^2 - 500 R^3 = 8.1

Hence R = .259 m and 0.397 m

42 ) The similar equation just changes the sign

K ( R - .5) = k Q^2/R^2

250 R^2 + 500 R^3 = 9*109 * 9*10-10

R = 0.553 m